# GUARD: Glocal Uncertainty-Aware Robust Decoding for Effective and Efficient Open-Ended Text Generation

### v1

@@ Acknowledgments
This work was partially supported by Henan Provincial Center for Outstanding Overseas Scientists (No.GZS2025004), MOE Liberal Arts and Social Sciences Foundation (No.23YJAZH210), Major Program of National Social Science Foundation (No.23&ZD309), High Level Talent International Training Program of Henan Province (No. GCC2025010), the Chinese Scholarship Council (Grant No.202308410339), Matthias Aßenmacher received funding from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) as part of BERD@NFDI, under grant number 460037581. Julian Rodemann acknowledges support by the Federal Statistical Office of Germany within the co-operation project "Machine Learning in Official Statistics" as well as by the Bavarian Institute for Digital Transformation (bidt) and the Bavarian Academy of Sciences (BAdW) within a graduate scholarship. We wish to thank the Munich Center for Machine Learning (MCML), and the Department of Statistics, LMU Munich. Henan Provincial Science and Technology Project (No. 252102210150), Key Scientific Research Project for Universities in Henan Province (No. 25A520014). Finally, we would like to thank all the anonymous reviewers and area chair(s) for the constructive discussion and for their suggestions throughout the review process.

### v2

@@ Acknowledgments
This work was partially supported by the MOE Liberal Arts and Social Sciences Foundation (No.23YJAZH210), Major Program of National Social Science Foundation (No.23&ZD309), Henan Provincial Center for Outstanding Overseas Scientists (No.GZS2025004), High Level Talent International Training Program of Henan Province (No.GCC2025010), Henan Provincial Science and Technology Project (No.252102210150), Key Scientific Research Project for Universities in Henan Province (No.25A520014), and the Chinese Scholarship Council (Grant No.202308410339). Moreover, Matthias Aßenmacher received funding from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) as part of BERD@NFDI, under Grant No.460037581. Julian Rodemann acknowledges the funding support from the Federal Statistical Office of Germany within the co-operation project "Machine Learning in Official Statistics", the Bavarian Institute for Digital Transformation (bidt) and the Bavarian Academy of Sciences (BAdW) within a graduate scholarship. We also thank the support from the Munich Center for Machine Learning (MCML), and the Department of Statistics, LMU Munich. Last but not least, we thank all the anonymous reviewers and area chair(s) for their constructive suggestions and help throughout the review process.